# Understanding Model Calibration - A gentle introduction and visual exploration of calibration and the expected calibration error (ECE)

### v4

@@ 2 Evaluating Calibration - Expected Calibration Error (ECE)
$\hat{y}_{i}$ is the model’s predicted class (arg max) for sample $i$ and $y_{i}$ is the true label for sample $i$. ${\displaystyle\bm{1}}$ is an indicator function, meaning when the predicted label $\hat{y}_{i}$ equals the true label $y_{i}$ it evaluates to 1, otherwise 0. Let’s look at an example, which will clarify acc, conf and the whole binning approach in a visual step-by-step manner.
@@ 2.1 ECE - Visual Step by Step Example
We can get the empirical probability of a sample falling into $B_{5}$, by assessing how many out of all $9$ samples fall into $B_{5}$, see ( $\mathbf{(\;1\;)}$. We then get the average accuracy for $B_{5}$, see ( $\mathbf{(\;2\;)}$ and lastly the average estimated probability for $B_{5}$, see ( $\mathbf{(\;3\;)}$. Repeat this for all bins and in our small example of $9$ samples we end up with an ECE of $0.10445$. A perfectly calibrated model would have an ECE of 0.

### v5

@@ 2 Evaluating Calibration - Expected Calibration Error (ECE)
$\hat{y}_{i}$ is the model’s predicted class (arg max) for sample $i$ and $y_{i}$ is the true label for sample $i$. $\bm{1}$} is an indicator function, meaning when the predicted label $\hat{y}_{i}$ equals the true label $y_{i}$ it evaluates to 1, otherwise 0. Let’s look at an example, which will clarify acc, conf and the whole binning approach in a visual step-by-step manner.
@@ 2.1 ECE - Visual Step by Step Example
We can get the empirical probability of a sample falling into $B_{5}$, by assessing how many out of all $9$ samples fall into $B_{5}$, see $\mathbf{(\;1\;)}$. We then get the average accuracy for $B_{5}$, see $\mathbf{(\;2\;)}$ and lastly the average estimated probability for $B_{5}$, see $\mathbf{(\;3\;)}$. Repeat this for all bins and in our small example of $9$ samples we end up with an ECE of $0.10445$. A perfectly calibrated model would have an ECE of 0.